OpenStudy (anonymous):

solve each inequality or equation for x 2|x+2|+4<28 Can Someone Please Help me

OpenStudy (anonymous):

subtract 4 get \[2|x+2|<24\] then divide by 2 and get \[|x+2|=12\]

OpenStudy (anonymous):

then you have to solve two equations, get two answers the equations are \[x+2=12\] of \[x+2=-12\]

OpenStudy (anonymous):

thank you